OpenStudy (anonymous):

What is the mean of the following data without the outlier? 41, 44, 50, 105 52 60 33.75 45

OpenStudy (anonymous):

the meaning (without the outlier) is nothing more than 6 integers and one decimal that all lie from 33.75-60. That's about it, unless you want to start taking averages and stuff like that

OpenStudy (anonymous):

I know that but whats the answer

OpenStudy (anonymous):

lol, I misread "mean" as meaning, so I was thinking of it in a similar way to "what's the meaning of life?". The mean is 41+44+50+52+60+33.75+45 all divided by 7 (remember to discard the outlier). 85+102+105+33.75)/7=202+90+33.75)/7=295.75/7..that's 40+2+1.75/7..42+1/7+3/28......42+7/28=42 1/4

OpenStudy (anonymous):

Please just give me the answer! :D

OpenStudy (anonymous):

42.25 should be it dude

OpenStudy (anonymous):

GRACIA! AMIGO!

OpenStudy (anonymous):

you are wrong